0.07g of anhydride of carbon occupies 56dm³ at stp when vaporized.and contains 14.29% by mass hydrogen.what is the mass of the hydrocarbon.


1
Expert's answer
2021-06-14T13:36:20-0400

Since the sample is a hydrocarbon, it contains Carbon and hydrogen in a specific ratio.

"1GMW=22.4dm^3"

Molar mass of Hydrocarbon "=56dm^3"

"1GMW" of "CO_2=12+2\times24=44gm=12gm Carbon"

"=" "12\over44" "\times 0.07=0.0191"

"\%" Carbon in the sample"=85.71\%"

"\%" Hydrogen in the sample "=14.29\%"

Empirical formualr"=" "85.71\over 12.01" "=7.14"

."14.29\over 1.0079" "=14.29"

"=1:2"

Formulae of Hydrocarbon"=(CH_2)_n"

"=(12.011+2\times 1.0079)\times n=56"

"=14.0268n=56"

"n=3.99=4"

"C_4H_8"

Mass of Hydrocarbon "=(12.011\times 4)+(1.0079\times 8)"

"=56.1072g"